QUESTION 9-4. Determine the angular acceleration of AB as CD moves with the angular speed and acceleration shown.

0.6 m ос 0.6 m B ( = 4 rads - 2 radís 0.3 m А DO
Answer #1

Given WBC Vc = 1.8m/s Tc/IC VBC BIC 0.6m (Ac) t = 3.6 m/s² WBC =0 12 BC (Ac)n = 3.6 mis? 0.6m (aolt (AB)n = 10.8 m/s²

IC is at o, so WBC =0 VB = vc - 0.9x2 = 1.8 mis 2 (ac) x0.9 = 3.6 m/s² V 4x0.9 =3.6 mis? (act (aon 1.82 - = 10.8 m/s?t 0.3 AB